# Directional Diffusion-Style Code Editing Pre-training

### v1

@@ IV-A Research Questions
To answer this question, we pre-train various ablated versions of DivoT5 by removing one pre-training task each, in order to evaluate the effect of each pre-training task.
@@ IV-B Pre-training datasets
CodeChangeNet [11] 35k+ 1.5M+
CommitPack-Java [30] 130k+ 3.7M+
@@
We use real code editing data to train DivoT5, using a code change dataset collected from the open-source repository Github, including CodeChangeNet [11] and CommitPack [30]. CodeChangeNet includes six popular programming languages (i.e., Go, Java, JavaScript, PHP, Python, and Ruby), over 35,000 projects, and over 1.5 million commit data. CommitPack comprises 4TB of GitHub commit data, covering 350 programming languages. To conserve computational resources, we select the Java programming language from CommitPack for our experiments. Specifically, the Java dataset includes over 165,000 projects and more than 3.7 million data instances. Generally, more data can lead to better model performance.
@@ IV-C Downstream Tasks
NL-guided Code Editing Code Review CodeReview 13,753 1,719 1,718
Bug Fixing Small 46,628 5,828 5,831
Medium 52,324 6,542 6,538
@@
NL-
CodeRefinement
150,406 13,103 13,104
Code-Only Editing Code Refinement Small 46,680 5,835 5,835
Medium 52,364 6,545 6,545
@@
Code editing tasks mainly involve the process of editing old code into new code. During the code editing process, there may be additional constraints that restrict the direction of evolution, such as adhering to natural language requirements. Therefore, the effectiveness of a code editing pre-trained model should be evaluated on two scenarios: the NL-guided code editing scenario and the code-only editing scenario. In particular, we focus on three tasks  (i.e., the automated code review task [3, 8], the bug fixing task [31], and the NL-based code refinement task [7]) for the first scenario and the code refinement task [6, 16] for the second scenario. We further use the code translation task [32, 16], which is a non-editing task, for evaluation. Since code translation differs much from our pre-training tasks, the code translation task may help evaluate the generalizability of DivoT5. Table II presents the statistical information of the datasets related to these tasks.
@@ IV-C1 NL-guided code editing scenario
The automated code review task aims to modify old code and generate new code based on the guidance provided in the form of review comments in natural language. We use the CodeReview dataset [3] with the Java programming language and adopt the same data settings as Zhang et al. [8]. The bug fixing task aims to fix the bug existing in the old code. We use the bug-fixing dataset with additional contextual information [31, 33], such as developers’ natural language guidance, to help the model fix bugs more accurately. The NL-based code refinement task aims to enhance the old code based on the guidance provided in natural language. We conduct experiments using a large-scale NL-based code refinement dataset [7] that includes nearly 180k data samples across nine programming languages.
@@ IV-C2 Code-only editing scenario
The code refinement task aims to convert defective code into correct one without natural language guidance. We use the Java dataset collected by Tufano et al. [6] to evaluate the model’s performance on this task. This dataset comprises two versions of benchmarks: the Refine-small version with code tokens fewer than 50, and the Refine-medium version with code tokens ranging between 50 and 100. We follow the settings of CodeXGLEU [16] and use the same metrics to evaluate our model.
@@ IV-C3 Non-editing scenario
The code translation task involves translating a program written in one programming language into another programming language. This task assists developers in code migration and cross-platform development using different programming languages. We use the CodeTrans dataset [32, 16] as our benchmark for this task, which includes Java and CSharp code samples, and evaluate it with the same metrics as the code refinement task.
@@ IV-D Metrics
We follow the existing research in the field of code generation [16] and code evolution [8] and select three metrics to evaluate the generated code: Exact Match, BLEU, and CodeBLEU.
@@ IV-D1 Exact Match (EM)
The EM score [34], often used in natural language processing tasks such as question answering, evaluates the accuracy of model predictions by measuring the proportion of examples for which the predicted answer exactly matches the ground truth answer.
@@ IV-D2 Bilingual Evaluation Understudy (BLEU)
The BLEU [35] score, also commonly used in natural language processing, evaluates the accuracy of model predictions by measuring the ratio of n-gram matches between the generated answer and the ground truth answer. In practice, the BLEU-4 score is often reported, with a higher BLEU-4 score indicating better quality of the generated answer.
@@ IV-D3 CodeBLEU (C-BLEU)
The CodeBLEU [36] score, adapted from BLEU and designed specifically for programming language processing, evaluates the accuracy of model predictions by summing four weighted parts. The equation for calculating CodeBLEU is as follows:
@@ IV-E Baselines
Despite the availability of numerous large-scale code pre-trained models[37, 38], fine-tuning is challenging due to their massive parameter sizes. Therefore, we compare the fine-tuned performance of models with similar parameter sizes and the few-shot performance of billion-scale models. Since our approach draws on the diffusion style to simulate the step-by-step editing process in real-world development scenarios, we focus our comparisons on auto-regressive code models rather than traditional diffusion models. By referencing diffusion concepts to generate data for training, our approach adapts these principles within an auto-regressive framework, making direct comparisons with traditional diffusion models (i.e., non-autoregressive framework) outside the scope of this work.
@@ IV-E1 Code pre-trained models
CodeT5 is a pre-trained model based on the encoder-decoder structure of T5 [20]. During the pre-training process, CodeT5 applies four pre-training tasks: masked span prediction, identifier tagging, masked identifier prediction, and bimodal dual generation. Identifier tagging is used to label whether a token is a code identifier, while the masked identifier prediction task is used to generate the masked identifier content. Experimental results demonstrate that CodeT5 can better capture code semantics and outperforms previous approaches on code-related tasks. In our evaluation, we use two variants of CodeT5: CodeT5-small (60M) and CodeT5-small (220M). Compared to CodeT5, CodeT5+ introduces more pre-training data and tasks, such as span denoising, contrastive learning, and causal LM. In our evaluation, we use the 220M parameter version of CodeT5+ (which is of the same scale as DivoT5-base) for comparison.
@@ IV-E2 Code editing pre-trained models
In our evaluation, we use four code editing pre-trained models: CodeReviewer, CCT5, CoditT5, and CodeEditor. Among the four models, CodeEdit has 60M parameters, and the other three have 220M parameters. CodeReviewer is a pre-trained model on a real-world dataset in code change scenarios. The primary pre-training task of CodeReviewer is to utilize a masked language modeling approach to mask the sequence representing line level code differences, and then predict the masked content. Additionally, CodeReviewer also incorporates a pre-training task of generating commit messages from code differences to enhance the model’s ability to generate text. CCT5 is a pre-trained model on a large dataset consisting of 1.5M+ pairwise data of code changes and commit messages from popular GitHub projects. In the pre-training process, CCT5 leverages the difference sequence of code and applies the SMLM task to learn the representation of code changes. For example, CCT5 first masks the information in the difference sequence at the granularity of code lines, then use other information from the difference and the commit message to predict the masked content. CoditT5 designs denoising pre-training tasks specifically for code editing. CoditT5 first calculates the probability of code edits in downstream tasks, including the probabilities of token insertion, deletion, and modification. Based on these probabilities, CoditT5 introduces noise to the old code. Finally, the model learns the code editing process by reconstructing the old code from the perturbed code. CodeEditor first disrupts the old code based on probabilities related to editing operations. Then, CodeEditor utilizes the existing generation model [16] to fill in the disrupted parts as input. Similar to CoditT5, CodeEditor also employs a denoising approach to recover the old code from the disrupted code.
IV-E3 Models with billion-scale parameters
@@
To more comprehensively evaluate the contribution of our model, we compared its performance with billion-scale parameter models, including ChatGPT (gpt-3.5-turbo) [37] and DeepSeekCoder-6.7B-instruct [38], Meta-Llama-3.1 (8B-instruct) [39], on these tasks. These models have been widely adopted due to their strong generative performance in general code generation. Given the large size of these models, fine-tuning is difficult due to limited resources, so we employ a few-shot approach. For the few-shot prompts, we refer to existing work [40], providing scenario descriptions of the problems and setting a low temperature to obtain the most stable and accurate results possible. We refine prompts several times based on  [40], with a temperature of 0 for reproducibility, and choose a top-performing few-shot example from the training set after 5 attempts. All the prompts used in our experiments can be found in [41]. Additionally, since these models are not fine-tuned on the training set, to ensure a fairer comparison, we remove the code style elements such as casing, line breaks, and comments from both the model-generated results and the ground truth in the test set. However, tasks related to code editing have strong domain-specific characteristics, which often limit the performance of these models.
@@ IV-F Training Details
In the pre-training phase, we employ the popular PyTorch [42] framework for model pre-training on the Linux Ubuntu operating system. Our pre-training process relies on 4 NVIDIA GeForce RTX 3090 GPUs, with each GPU having a memory capacity of 24GB. Additionally, our system has a memory capacity of 251GB. In the small version of pre-training on CodeT5 with 60M parameters, we use a learning rate of 3e-4. For the base version of pre-training on CodeT5 with 220M parameters, we use a learning rate of 2e-4. To expedite the model’s pre-training process, we employ the Accelerate [43] and DeepSpeed [44] frameworks with mixed-precision BF16 to improve the training speed.
During the fine-tuning process for downstream tasks, we maintain the same system environment and acceleration settings as in the pre-training phase. We set length restrictions for the inputs and outputs of each task based on statistical analysis of the dataset samples and existing works [8, 16, 7]. All models use the same length settings for the same dataset.
@@ V-A RQ1: Effectiveness of DivoT5
We evaluate the performance of DivoT5 on two types of code editing scenarios (i.e. NL-guided code editing scenario and code-only editing scenario) across four specific datasets, comparing it with existing code pre-trained models. If the model has already been evaluated on the datasets, we directly use the performance metrics provided in the original paper. Otherwise, we load the released model from Huggingface [45] and use the published code to evaluate it on the corresponding datasets. If the model mentioned in the original paper is not released yet, we only compare it with our experimental results on the same dataset that we use and indicate this with a ‘-’ in the table.
@@ V-A1 NL-guided code editing scenario
CodeReview BugFix NL-based Code Refinement
Models Small Medium
Metrics EM BLEU EM BLEU EM BLEU EM BLEU
CodeT5-small (60M) 32.25 85.86 33.10 73.97 25.45 85.01 28.07 78.06
CodeT5-base (220M) 34.46 86.21 36.84 75.45 30.47 85.68 29.44 78.64
CodeT5-large (770M) 34.63 86.10 36.55 75.50 29.83 85.86 29.81 79.17
CodeT5+ (220M) 32.36 85.99 34.27 73.88 25.19 84.66 30.24 78.64
CodeEditor (60M) 29.40 - - - - - - -
CoditT5 (220M) 37.19 - 37.52 - 29.96 - 28.95 78.98
CCT5 (220M) 36.96 87.65 36.72 74.95 30.15 84.78 29.66 78.55
CodeReviewer (220M) 37.14 87.30 36.36 74.28 31.16 85.34 30.67 78.96
DeepSeekCoder-FS (6.7B-instruct) 26.83 76.17 2.68 43.76 2.26 59.63 15.78 70.04
Meta-Llama-3.1-FS (8B-instruct) 30.15 71.39 2.30 34.02 2.26 47.54 18.28 65.45
ChatGPT-FS (gpt-3.5-turbo) 34.81 82.05 2.42 54.98 1.71 65.71 21.99 80.47
DivoT5-small (60M) 40.16 87.81 35.84 74.80 28.72 84.65 29.42 78.59
DivoT5-base (220M) 44.41 88.54 39.08 75.86 33.65 85.24 32.75 79.19
@@
Table III presents a comparison of the performance for the NL-guided code editing scenario, emphasizing the effectiveness of DivoT5. It can be observed that the DivoT5-base model demonstrates significant improvements across most metrics on all tasks compared to all other models, including 60M, 220M, 770M, and billion-scale parameters.
@@
For the automated code review task, the DivoT5-base model shows nearly a 29% improvement in exact match (EM) accuracy compared to the CodeT5-base model, with EM increasing from 34.46% to 44.41%. The DivoT5-small model surpasses even models with 220M parameters in all metrics. DivoT5-base also outperforms CoditT5 and CodeReviewer by over 18% in comparison to other code editing pre-trained models. Figure 4 depicts a case study showcasing different models’ outputs on the CodeReview dataset. To save space, we only compare the outputs of models with a parameter count of 220M and display the differences between the new code and the old code. In Figure 4, the old code’s functionality checks for null in the return value. The review comment suggests using “ReviewCategoryStrategy.NONE” instead of returning a null value. Specifically, the reviewer recommends replacing the return value in the last line of the old code but not changing the null value used for checks. The CodeT5 model returns the “ReviewCategoryStrategy.NONE” value, incorrectly removes the necessary condition check. CoditT5 and CodeT5+ mistakenly delete the original null check, while the CodeReviewer model outputs the old code without making any modifications. In this example, DivoT5 is able to generate the evolved new code correctly. This could be attributed to the design of DivoT5’s pre-training task, which enables the model to perceive which parts do not require modification (i.e., the $KSM_{ED}$ pre-training task).
@@
Similarly, for the bug fixing task, DivoT5 exceeds other models of same size in almost all metrics. While DivoT5-base shows a slight decrease in BLEU score compared to CodeT5-base in the medium version of the dataset, it increases EM by over 10%. Given that the EM metric is typically more crucial than the BLEU score in indicating correct bug fixes, this improvement is significant. In the large-scale, multi-language code refinement dataset, the DivoT5-base continues to enhance the EM metric. Compared to the similarly sized CodeT5-base model, EM increases from 29.44% to 32.75%, an improvement of over 11.2%. Since the fine-tuned weights of CodeT5-base from [7] are unavailable, we re-fine-tuned CodeT5 on this dataset for 13 epochs over 44 hours and 52 minutes until performance plateaued on the validation set. Notably, despite being trained only on Java data from the CommitPack dataset due to resource constraints, DivoT5 achieves a higher EM score than the CodeReviewer model, which was trained on data from multiple programming languages.
When compared to billion-scale models in few-shot settings, we observe that these models generally show limited performance in NL-guided code editing scenarios. This is because code editing tasks are highly domain-specific, and the few-shot approach often struggles to learn sufficient domain knowledge. This observation aligns with existing research findings [40]. Additionally, in tasks such as automated code review and NL-based code refinement, the natural language descriptions are more closely related to the actual code changes, resulting in better performance from few-shot models. However, for bug-fixing tasks, the natural language descriptions typically provide relevant context rather than directly corresponding to the actual changes, leading to poorer performance from these models.

### v2

@@ IV-A Research Questions
To answer this question, we pre-train various ablated versions of DivoT5 by removing one pre-training task each, in order to evaluate the effect of each pre-training task. We conduct these ablation studies on the 60M-parameter variant of DivoT5 to enable efficient experimentation under limited computational resources, following a common practice in prior works [guo2024deepseekcoder, zhu2024grammart5].
@@ IV-B Pre-training datasets
CodeChangeNet [lin2023cct5] 35k+ 1.5M+
CommitPack-Java [commitpack_muennighoff2023octopack] 130k+ 3.7M+
@@
We use real code editing data to train DivoT5, using a code change dataset collected from the open-source repository Github, including CodeChangeNet [lin2023cct5] and CommitPack [commitpack_muennighoff2023octopack]. CodeChangeNet includes six popular programming languages (i.e., Go, Java, JavaScript, PHP, Python, and Ruby), over 35,000 projects, and over 1.5 million commit data. CommitPack comprises 4TB of GitHub commit data, covering 350 programming languages. To conserve computational resources, we select the Java programming language from CommitPack for our experiments. Specifically, the Java dataset includes over 165,000 projects and more than 3.7 million data instances. Generally, more data can lead to better model performance.
@@ IV-C Downstream Tasks
NL-guided
Code Editing
@@
CodeReview 13,753 1,719 1,718
NL-CodeRefinement 150,406 13,103 13,104
Bug Fixing BugFixing-S 46,628 5,828 5,831
BugFixing-M 52,324 6,542 6,538
Code-Only
Editing
Code Refinement Refine-S 46,680 5,835 5,835
Refine-M 52,364 6,545 6,545
CodeReview-CodeOnly 13,753 1,719 1,718
@@
Code editing tasks mainly involve the process of editing old code into new code. During the code editing process, there may be additional constraints that restrict the direction of evolution, such as adhering to natural language requirements. Therefore, the effectiveness of a code editing pre-trained model should be evaluated on two scenarios: the NL-guided code editing scenario and the code-only editing scenario. In particular, we focus on two tasks  (i.e., the NL-based code refinement task [tufano2021codereview, coditt52022, codereviewer] and the bug fixing task [tufano2019learning_bugfixing]) for the first scenario and the code refinement task [tufano2019coderefine, codexglue2021] for the second scenario. We further use the code translation task [chen2018codetrans, codexglue2021], which is a non-editing task, for evaluation. Since code translation differs much from our pre-training tasks, the code translation task may help evaluate the generalizability of DivoT5. Table II presents the statistical information of the datasets related to these tasks.
@@ IV-C1 NL-guided code editing scenario
The NL-based code refinement task focuses on improving existing code based on guidance provided in natural language. We first conduct experiments on the CodeReview dataset [tufano2021codereview], which is written in Java, following the same data setup as Zhang et al.[coditt52022]. To further evaluate the model’s capability, we also utilize a large-scale multilingual dataset[codereviewer], which contains nearly 180k samples spanning nine programming languages. The bug fixing task aims to fix the bug existing in the old code. We use the bug-fixing dataset with additional information [tufano2019learning_bugfixing, chakraborty2021multi_editing], such as developers’ natural language guidance, to help the model fix bugs more accurately.
@@ IV-C2 Code-only editing scenario
The code refinement task aims to convert defective code into correct one without natural language guidance. We use the Java dataset collected by Tufano et al. [tufano2019coderefine] to evaluate the model’s performance on this task. This dataset comprises two versions of benchmarks: the Refine-small version with code tokens fewer than 50, and the Refine-medium version with code tokens ranging between 50 and 100. We follow the settings of CodeXGLEU [codexglue2021] and use the same metrics to evaluate our model. To more comprehensively evaluate this scenario, we construct a variant of the CodeReview [tufano2021codereview] dataset, referred to as CodeReview-CodeOnly, by removing the natural language input. This version retains the same number of samples as the original dataset but omits the NL comments from the input, enabling us to isolate and assess the model’s performance in the absence of natural language guidance.
@@ IV-C3 Non-editing scenario
The code translation task involves translating a program written in one programming language into another programming language. This task assists developers in code migration and cross-platform development using different programming languages. We use the CodeTrans dataset [chen2018codetrans, codexglue2021] as our benchmark for this task, which includes Java and CSharp code samples, and evaluate it with the same metrics as the code refinement task.
@@ IV-D Metrics
We follow the existing research in the field of code generation [codexglue2021] and code evolution [coditt52022] and select four metrics to evaluate the generated code: Exact Match, BLEU, CodeBLEU, and SARI.
@@ IV-D1 Exact Match (EM)
The EM score [Xinyun2018EM], often used in natural language processing tasks such as question answering, evaluates the accuracy of model predictions by measuring the proportion of examples for which the predicted answer exactly matches the ground truth answer.
@@ IV-D2 Bilingual Evaluation Understudy (BLEU)
The BLEU [Kishore2002BLEU] score, also commonly used in natural language processing, evaluates the accuracy of model predictions by measuring the ratio of n-gram matches between the generated answer and the ground truth answer. In practice, the BLEU-4 score is often reported, with a higher BLEU-4 score indicating better quality of the generated answer.
@@ IV-D3 CodeBLEU (C-BLEU)
The CodeBLEU [Shuo2020CodeBLEU] score, adapted from BLEU and designed specifically for programming language processing, evaluates the accuracy of model predictions by summing four weighted parts. The equation for calculating CodeBLEU is as follows:
@@
IV-D4 System output Against References and Input (SARI)
Evaluating code editing models requires not only output similarity but also an accurate assessment of how well the model performs actual edit operations [editprogress]. The SARI [sari] metric, originally proposed for text simplification, evaluates the quality of edits by comparing the system output against both the input and the reference. Specifically, it separately measures the precision and recall of three types of operations: additions, deletions, and keeps. Unlike similarity-based metrics such as BLEU or CodeBLEU, SARI explicitly focuses on the correctness of edit operations, making it more suitable for code editing tasks.
@@ IV-E Baselines
Despite the availability of numerous large-scale code pre-trained models [chatgpt2024, guo2024deepseekcoder], fine-tuning is challenging due to their massive parameter sizes. Therefore, we compare the fine-tuned performance of models with similar and larger parameter sizes and the few-shot performance of billion-scale instruct models. Since our approach draws on the diffusion style to simulate the step-by-step editing process in real-world development scenarios, we focus our comparisons on auto-regressive code models rather than traditional diffusion models. By referencing diffusion concepts to generate data for training, our approach adapts these principles within an auto-regressive framework, making direct comparisons with traditional diffusion models (i.e., non-autoregressive framework) outside the scope of this work.
@@ IV-E1 Code pre-trained models
CodeT5 is a pre-trained model based on the encoder-decoder structure of T5 [t52020]. During the pre-training process, CodeT5 applies four pre-training tasks: masked span prediction, identifier tagging, masked identifier prediction, and bimodal dual generation. Identifier tagging is used to label whether a token is a code identifier, while the masked identifier prediction task is used to generate the masked identifier content. Experimental results demonstrate that CodeT5 can better capture code semantics and outperforms previous approaches on code-related tasks. In our evaluation, we use three variants of CodeT5: CodeT5-small (60M), CodeT5-base (220M), and CodeT5-large (770M). Compared to CodeT5, CodeT5+ introduces more pre-training data and tasks, such as span denoising, contrastive learning, and causal LM. In our evaluation, we use the 220M parameter version of CodeT5+ (which is of the same scale as DivoT5-base) for comparison. CodeLlama [codellama] is a family of decoder-only models pre-trained on large-scale corpora of both natural language and code, with variants specifically optimized for code understanding and generation. It has shown strong performance across a wide range of code-related benchmarks. In our experiments, we use the 6.7B base version of CodeLlama as a representative and competitive LLM baseline to compare against DivoT5, highlighting the advantages of our task-specific pretraining approach.
@@ IV-E2 Code editing pre-trained models
In our evaluation, we use four code editing pre-trained models: CodeReviewer, CCT5, CoditT5, and CodeEditor. Among the four models, CodeEdit has 60M parameters, and the other three have 220M parameters. CodeReviewer is a pre-trained model on a real-world dataset in code change scenarios. The primary pre-training task of CodeReviewer is to utilize a masked language modeling approach to mask the sequence representing line level code differences, and then predict the masked content. Additionally, CodeReviewer also incorporates a pre-training task of generating commit messages from code differences to enhance the model’s ability to generate text. CCT5 is a pre-trained model on a large dataset consisting of 1.5M+ pairwise data of code changes and commit messages from popular GitHub projects. In the pre-training process, CCT5 leverages the difference sequence of code and applies the SMLM task to learn the representation of code changes. For example, CCT5 first masks the information in the difference sequence at the granularity of code lines, then use other information from the difference and the commit message to predict the masked content. CoditT5 designs denoising pre-training tasks specifically for code editing. CoditT5 first calculates the probability of code edits in downstream tasks, including the probabilities of token insertion, deletion, and modification. Based on these probabilities, CoditT5 introduces noise to the old code. Finally, the model learns the code editing process by reconstructing the old code from the perturbed code. CodeEditor first disrupts the old code based on probabilities related to editing operations. Then, CodeEditor utilizes the existing generation model [codexglue2021] to fill in the disrupted parts as input. Similar to CoditT5, CodeEditor also employs a denoising approach to recover the old code from the disrupted code.
IV-E3 Instruct-Models with billion-scale parameters
@@
To more comprehensively evaluate the effectiveness of our model, we compare it against several billion-scale instruction-tuned LLMs that are widely used in code generation and editing tasks [chatgpt2024, guo2024deepseekcoder, llama3]. These include ChatGPT (gpt-3.5-turbo) [chatgpt2024], DeepSeekCoder-6.7B-Instruct [guo2024deepseekcoder], and Meta-Llama-3.1 (8B-Instruct) [llama3]. These models are known for their strong general-purpose code generation capabilities and are representative of current state-of-the-art instruction-tuned LLMs. Given the high computational cost of fine-tuning such large models on our task, we evaluate them in a few-shot setting instead. Following prior work [chatgpt_refine], we design prompts that include scenario-specific descriptions of the editing task and use a low temperature to ensure consistent and accurate outputs. Specifically, we refine prompts several times based on  [chatgpt_refine], with a temperature of 0 for reproducibility, and choose a top-performing few-shot example from the training set after 5 attempts. All the prompts used in our experiments can be found in [divot5]. Additionally, since these models are not fine-tuned on the training set, to ensure a fairer comparison, we remove the code style elements such as casing, line breaks, and comments from both the model-generated results and the ground truth in the test set. However, tasks related to code editing have strong domain-specific characteristics, which often limit the performance of these models.
@@ IV-F Training Details
In the pre-training phase, we employ the popular PyTorch [pytorch] framework for model pre-training on the Linux Ubuntu operating system. Our pre-training process relies on 4 NVIDIA GeForce RTX 3090 GPUs, with each GPU having a memory capacity of 24GB. Additionally, our system has a memory capacity of 251GB. In the small version of pre-training on CodeT5 with 60M parameters, we use a learning rate of 3e-4. For the base version of pre-training on CodeT5 with 220M parameters, we use a learning rate of 2e-4. To expedite the model’s pre-training process, we employ the Accelerate [acclerate] and DeepSpeed [deepspeed] frameworks with mixed-precision BF16 to improve the training speed.
During the fine-tuning process for downstream tasks, we maintain the same system environment and acceleration settings as in the pre-training phase. We set length restrictions for the inputs and outputs of each task based on a statistical analysis of the dataset samples and existing works [coditt52022, codexglue2021, codereviewer]. All models use the same length settings for the same dataset.
@@ V-A RQ1: Effectiveness of DivoT5
We evaluate the performance of DivoT5 on two types of code editing scenarios (i.e., NL-guided code editing scenario and code-only editing scenario) across seven specific datasets, comparing it with existing code pre-trained models. If the model has already been evaluated on the target datasets, we directly report the performance metrics from the original paper. Otherwise, we download the released model from HuggingFace [huggingface] and use the official code to evaluate it on the corresponding datasets. For models whose checkpoints or code are not publicly available, we only include a comparison with our results on the same dataset, and mark missing entries with a ‘-’ in the table. We use the symbol α in the result tables to indicate models whose performance is not statistically significantly different from DivoT5, with both DivoT5 and the corresponding baselines jointly marked to reflect their comparable results.
@@ V-A1 NL-guided code editing scenario
Models CodeReview
NL-
CodeRefinement
BugFixing-S BugFixing-M
Metrics EM BLEU SARI EM BLEU SARI EM BLEU SARI EM BLEU SARI
CodeT5-small (60M) 32.25 85.86 69.79 28.07 78.06 69.82 33.10 73.97 58.45 25.45 85.01α 55.62
CodeT5-base (220M) 34.46 86.21 70.94 29.44 78.64 70.61 36.84 75.45 59.85 30.47 85.68α 57.10
CodeT5-large (770M) 34.63 86.10 71.28 29.81 79.17α 71.07 36.55 75.50α 59.56 29.83 85.86 57.03
CodeT5+ (220M) 32.36 85.99 69.71 30.24 78.64 71.07 34.27 73.88 58.87 25.19 84.66 55.22
CodeEditor (60M) 29.40 - - - - - - - - - - -
CoditT5 (220M) 37.19 86.58 70.33 28.95 78.98α 70.62 37.52 75.39 59.72 29.96 85.75 56.35
CCT5 (220M) 36.96 87.65 70.25 29.66 78.55 70.48 36.72 74.95 59.64 30.15 84.78 57.02
CodeReviewer (220M) 37.14 87.30 71.21 30.67 78.96α 71.49 36.36 74.28 59.57 31.16 85.34α 57.05
CodeLlama (6.7B-base) 31.49 85.19 67.44 31.65 79.07α 68.64 40.44 74.36 60.94α 29.40 83.91 56.62
DeepSeekCoder-FS (6.7B-instruct) 26.83 76.17 66.26 15.78 70.04 60.49 2.68 43.76 47.59 2.26 59.63 46.31
Meta-Llama-3.1-FS (8B-instruct) 30.15 71.39 66.91 18.28 65.45 61.59 2.30 34.02 47.25 2.26 47.54 46.14
ChatGPT-FS (gpt-3.5-turbo) 34.81 82.05 68.39 21.99 80.47 65.24 2.42 54.98 48.07 1.71 65.71 47.43
DivoT5-small (60M) 40.16 87.81 72.00 29.42 78.59 70.93 35.84 74.80 59.40 28.72 84.65 56.78
DivoT5-base (220M) 44.41 88.54 73.51 32.75 79.19α 72.09 39.08 75.86α 60.67α 33.65 85.24α 58.12
@@
Table III presents a comparison of the performance for the NL-guided code editing scenario, emphasizing the effectiveness of DivoT5. It can be observed that the DivoT5-base model demonstrates significant improvements across most datasets compared to all other models, including 60M, 220M, 770M, and billion-scale parameters.
For the NL-based code refinement task, DivoT5-base achieves a substantial improvement over CodeT5-base, with a 29% relative gain in the exact match metric (from 34.46 to 44.41), and consistently stronger results across BLEU and SARI. Notably, the smaller DivoT5-small model even surpasses models with 220M parameters in all evaluation metrics, demonstrating the effectiveness of our pretraining objectives even at a smaller scale. Compared to other pre-trained editing models such as CoditT5 and CodeReviewer, DivoT5-base also achieves over 18% higher performance in the exact match metric, further validating its superior ability to model real-world code edits. In the large-scale, multilingual NL-CodeRefinement dataset, which spans nine programming languages, DivoT5-base continues to outperform strong baselines. It achieves 32.75 in exact match and 72.09 in SARI, both improvements over CodeT5-base, which reaches 29.44 and 70.61 respectively. These results highlight DivoT5’s robustness and generalization capability across diverse codebases and natural language annotations. Since the fine-tuned weights of CodeT5-base from [codereviewer] are unavailable, we re-fine-tuned CodeT5 on this dataset for 13 epochs over 44 hours and 52 minutes until performance plateaued on the validation set.
@@
Figure 4 depicts a case study showcasing different models’ outputs on the CodeReview dataset. To save space, we only compare the outputs of models with a parameter count of 220M and display the differences between the new code and the old code. In the Figure 4, the old code’s functionality checks for null in the return value. The review comment suggests using “ReviewCategoryStrategy.NONE” instead of returning a null value. Specifically, the reviewer recommends replacing the return value in the last line of the old code but not changing the null value used for checks. The CodeT5 model returns the “ReviewCategoryStrategy.NONE” value, incorrectly removes the necessary condition check. CoditT5 and CodeT5+ mistakenly delete the original null check, while the CodeReviewer model outputs the old code without making any modifications. In this example, DivoT5 is able to generate the evolved new code correctly. This could be attributed to the design of DivoT5’s pre-training task, which enables the model to perceive which parts do not require modification (i.e., the $KSM_{ED}$ pre-training task). Notably, despite being trained only on Java data from the CommitPack dataset due to resource constraints, DivoT5 achieves a higher EM score than the CodeReviewer model, which was trained on data from multiple programming languages.
@@
Similarly, for the bug fixing task, DivoT5 outperforms all models of comparable size across most evaluation metrics. On the BugFixing-S dataset, CodeLlama (6.7B-base) achieves the highest scores with 40.44 in exact match and 60.94 in SARI, demonstrating strong performance on short, localized edits. However, DivoT5-base closely follows with 39.08 in exact match and 60.67 in SARI, despite having over 30× fewer parameters. This result underscores the effectiveness of our fine-grained edit-oriented pretraining in achieving competitive performance without relying on massive model scale. On the more challenging BugFixing-M dataset, which involves longer and more complex edits, DivoT5-base further strengthens its lead with 33.65 in exact match and 58.12 in SARI, outperforming all other models, including those with significantly larger parameter counts. While DivoT5-base shows a slight decrease in BLEU score compared to CodeT5-base in the BugFixing-M dataset, it increases EM by over 10%. Given that the EM metric is typically more crucial than the BLEU score in indicating correct bug fixes, this improvement is significant.
When compared to billion-scale models in few-shot settings, we observe that these models generally show limited performance in NL-guided code editing scenarios. This is because code editing tasks are highly domain-specific, and the few-shot approach often struggles to learn sufficient domain knowledge. This observation aligns with existing research findings [chatgpt_refine]. Additionally, in NL-based code refinement, the natural language descriptions are more closely related to the actual code changes, resulting in better performance from few-shot models. However, for bug-fixing tasks, the natural language descriptions typically provide relevant context rather than directly corresponding to the actual changes, leading to poorer performance from these models.